OpenStudy (anonymous):

Let y=3cosx −3. what is period and phase shift

OpenStudy (anonymous):

is this \[3\cos(x)-3\] which has period \[2\pi\] and on phase shift, or \[3\cos(x-3)\] which also has period \[2\pi\] and shifted 3 unit right

OpenStudy (anonymous):

Let y=−3sin[5(x−π/6)]+5. can you help me with this one also. I need the period and range

OpenStudy (anonymous):

Let y=sin[x+π/3]. i would greatly appreciate it if you can help me find the range for this

OpenStudy (anonymous):

range of sine is [-1,1]

OpenStudy (anonymous):

so range of \[-3\sin(x)\] would be [-3,3]